# Near Optimal Locality Sensitive Orderings in Euclidean Space

### v2

@@ Abstract
For a parameter $\varepsilon\in(0,1)$, a set of $\varepsilon$-locality-sensitive orderings (LSOs) has the property that for any two points, $p,q\in[0,1)^{d}$, there exist an order in the set such that all the points between $p$ and $q$ (in the order) are $\varepsilon$-close to either $p$ or $q$. Since the original construction of LSOs can not be (significantly) improved, we present a construction of modified LSOs, that yields a smaller set, while preserving their usefulness. Specifically, the resulting set of LSOs has size $M=O(\mathcal{E}^{d-1}\log\mathcal{E})$, where $\mathcal{E}=1/\varepsilon$. This improves over previous work by a factor of $\mathcal{E}$, and is optimal up to a factor of $\log\mathcal{E}$.
As a consequence we get a flotilla of improved dynamic geometric algorithms, such as maintaining bichromatic closest pair, and spanners, among others. In particular, for geometric dynamic spanners the new result matches (up to the aforementioned $\log\mathcal{E}$ factor) the lower bound, Specifically, this is a near-optimal simple dynamic data-structure for maintaining spanners under insertions and deletions.
@@ 1 Introduction
Chan et al. [CHJ20] showed that there is a “small” set $\Pi$ of orderings, such that for all points $p,q\in\mathcal{H}$, there is an order $\sigma\in\Pi$ that is $\varepsilon$-local, where $\varepsilon\in(0,1)$ is a fixed parameter.
@@
For a pair of points $p,q\in\mathcal{H}=[0,1)^{d}$, an order $\sigma$ over the points of $\mathcal{H}$ is $\varepsilon$-local if
$\sigma(p,q)\quad\subseteq\quad\mathcalb{b}\left({p,\varepsilon\ell}\right)\cup\mathcalb{b}\left({q,\varepsilon\ell}\right),\qquad\text{where}\qquad\ell=\left\|{pq}\right\|,$
where $\mathcalb{b}\left({p,r}\right)$ denotes the ball of radius $r$ centered at $p$, see Figure 1.1 (i).
Namely, all the points between $p$ and $q$ in $\sigma$ are in the vicinities of $p$ and $q$ in $\mathcal{H}$. Informally, there is an order in $\Pi$ that respects the localities of $p$ and $q$. Chan et al., inspired by locality sensitive hashing, refereed to $\sigma$ as being a locality sensitive ordering (LSO).
Surprisingly, Chan et al. showed that one can compute such a “universal” set of orderings $\Pi$, of size $O(\mathcal{E}^{d}\log\mathcal{E})$, where $\mathcal{E}=1/\varepsilon$. This set of orderings can be easily computed, and computing the order of two points according to a specific order in the set can be done quickly. Having such a set of orderings, reduces some dynamic geometric problems to the maintenance of a “few” sorted lists of points under insertions and deletions. Thus, certain problems in $d$ dimensions, are reduced to a collection of problems in one dimension. This can be interpreted as a somewhat strange dimension reduction result, (multi)embedding $[0,1)^{d}$ in the real line, such that for any two points in the hypercube, there is at least one embedding that has certain desired geometric property.
For example, using these orderings one can easily maintain a spanner under insertion/deletion operations with logarithmic time per operation (ignoring terms depending on $\varepsilon$). Specifically, one maintains (under updates) sorted lists of the point set, according to each of the orderings in $\Pi$. Here, an edge in the spanner is present only if the two points are adjacent in one of these ordered lists. Note, that for each of these one dimensional instances, one maintains a dynamic ordered map of the points (e.g., balanced binary search tree).
The challenge
So to reduce the number of orderings further, we have to relax the requirement somehow. A natural first step is to enlarge the vicinity to $p$ and $q$ to be an “$\varepsilon$-hippodrome” region. Specifically, we require that
$\sigma(p,q)\quad\subseteq\quad\mathcalb{h}\left({pq,\varepsilon\ell}\right)=\mathcal{CH}\left({\bigl{.}\mathcalb{b}\left({p,\varepsilon\ell}\right)\cup\mathcalb{b}\left({q,\varepsilon\ell}\right)}\right),$
see Figure 1.1 (ii), where $\mathcalb{h}\left({pq,r}\right)$ is the $r$-hippodrome of $pq$. This condition is sufficient for some applications, but others require a stronger property of having a “gap” in the ordering.
Definition 1.2.
For $\gamma\in(0,1/4)$, with $\gamma\geq\varepsilon$ (e.g., $\gamma=1/4$), consider the requirement that
@@
$\sigma(p,q)\quad\subseteq\quad\mathcalb{h}\left({pq,\varepsilon\ell}\right)\cap\left({\Bigl{.}\mathcalb{b}\left({p,\gamma\ell}\right)\cup\mathcalb{b}\left({q,\gamma\ell}\right)}\right),\qquad\text{where}\qquad\ell=\left\|{pq}\right\|,$
see Figure 1.1 (iii). An order $\sigma$ that has the last property is an $(\varepsilon,\gamma)$-local for $p$ and $q$.
Computation model
@@
The model of computation used here is a unit-cost real RAM, supporting standard arithmetic operations and comparisons (but no floor function), augmented with standard bitwise-logical operations (bitwise-exclusive-or and bitwise-and), which are widely available as assembly commands on all modern general purpose CPUs, and programming languages. This computational model is reasonable, and is used (for example) in working with compressed quadtrees efficiently [Har11].
Previous related work
@@
LSOs were introduced by Chan et al. [CHJ20]. Variations of LSOs for doubling metrics, and other metric spaces, were studied by Filtser and Le [FL22] (Triangle LSO, Left-sided LSO), and Filtser [Fil23] (Triangle LSO, Rooted LSO). Since these works study LSO in the non-euclidean settings, they are not directly related to our work here (and they yield, naturally, worse bounds). LSOs were used to construct reliable spanners by Buchin et al. [BHO20, BHO22] in a surprisingly simple plug & play fashion. There is some work on quadtrees for hyperbolic space by Kisfaludi-Bak and van Wordragen [KW23], which potentially can lead for LSOs for such spaces.
@@
In this work, we show that there is a universal set of LSOs of $[0,1)^{d}$ that is of size $O(\mathcal{E}^{d-1}\log\mathcal{E})$, such that these orderings have the somewhat weaker $(\varepsilon,\gamma)$-locality property, where $\gamma\geq\varepsilon$ is a large fixed constant. This improves the result of Chan et al. by a factor of $1/\varepsilon$. Furthermore, for many of the applications using LSOs, one can replace the LSOs by the new set of “weaker” orderings, thus strictly improving their dependence on $\varepsilon$. In particular, for spanners the new construction enables maintaining dynamic spanners. Notably, up to $\log\mathcal{E}$ factor, the new construction matches the known lower bound on the number of edges in a spanner, in the static case, thus implying that the new orderings are almost optimal, and no approach based on LSOs can do (significantly) better in the Euclidean case.
Applications
Since we have to slightly weaken the definition of LSOs to get the improved size, we have to rederive the proofs of correctness for some of the applications. LSOs are especially useful in applications where one wants to solve some geometric problem under insertions and deletions of points. Since LSOs decouple the geometric problem from the data-structure problem (i.e., all one needs to do is to maintain sorted lists of points under insertions and deletions), this results in surprisingly simple data-structures that supports updates:
@@
    Bichromatic closest pair. In Section 4.2, we show how to use LSOs to dynamically maintain the bichromatic closest pair, as the shortest bichromatic edge in the locality graph is the desired approximation. This also readily leads to a data-structure that supports $(1+\varepsilon)$-approximate nearest-neighbor queries. We do not present the details for this here, as this result is, relatively speaking, less interesting.
@@
All these results improve by a factor of $\mathcal{E}$ over the previous results of Chan et al.. The new LSOs also improves, by a factor of $\mathcal{E}$, in a plug and play fashion other results, but the improvement is somewhat less interesting in these cases. This includes reliable spanners [BHO20], dynamic approximate minimum spanning trees [CHJ20], and vertex/edge fault-tolerant spanners [CHJ20].
Sketch of the old construction
We describe shortly the construction of LSOs by Chan et al. [CHJ20], see Section 2.1 for more details. Chat et al. reduced the problem of computing LSOs to the following problem – consider the integer grid $B=\left\llbracket 4\mathcal{E}\right\rrbracket^{d}$, where $\mathcal{E}=1/\varepsilon$ and $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$. An edge of $G$ is long if it corresponds to a pair of points of $B$ that are in distance at least (say) $2/\varepsilon$ from each other. The long edges corresponds to well-separated pairs of points, and these are the ones we must “serve”. (Understanding why we concentrate on the long edges requires a bit deeper dive into the details, see Section 2.1.) To this end, consider the complete graph $G$ over $B$. Constructing LSOs is then achieved by constructing LSOs for $B$. A single ordering (i.e., LSO) for $B$ is simply a Hamiltonian path of $G$, and the task at hand is to compute a small number of Hamiltonian paths of $G$ that visits all the long edges. Specifically, a Hamiltonian path (in a graph) induces an ordering of the vertices, thus Hamiltonian paths serves as a way to encode the desired orderings. Having a long edge in such a path, ensures this pair is consecutive in the associated ordering.
Walecki [Als08] showed a decomposition of the clique into Hamiltonian paths, and this leads to the desired LSOs of $B$. The problem with this approach is that it requires $\Omega(\mathcal{E}^{d})$ paths/LSOs. There is an additional blowup by a factor of $O(\log\mathcal{E})$ in the number of LSOs, when converting the LSOs of $B$ to LSOs of $[0,1)^{d}$.
Sketch of new approach
@@
Why does the size of LSOs matter? LSOs seems to be algorithmically equivalent to other geometric tools, such as well-separated pairs decomposition [CK95]. Thus, they seem to be a fundamental property of the underlying space. It is thus natural to quantify exactly what the minimum size of LSOs needed for $\mathbb{R}^{d}$, as a function of $d$ and $\varepsilon$. Our work thus can be interpreted as (almost) settling this question.
@@
We describe the old construction of Chan et al. in Section 2.1. We present some necessary standard results on nets/packings in Section 2.2. In Section 3.1 we describing how to construct the needed packings on projected points, and how to decompose such a packing into a packing of larger radius. At last, we are ready. We present, in Section 3.2, the construction of the orderings with the hippodrome property, and, in Section 3.3, we present the modified construction with the gap property.
@@ 2.1 Locality sensitive orderings
Our starting point is the result of Chan et al. [CHJ20]. Since we use the same framework, we quickly sketch the above construction (reproducing some relevant definitions).
@@
By fixing the root of the $\varepsilon$-quadtree to be $[0,2^{i})^{d}$, for $i=1,\ldots,\lambda$, one gets the following.
@@
Let $\mathcal{T}$ be a regular quadtree over $[0,2)^{d}$. There are $\varepsilon$-quadtrees $\mathcal{T}_{\varepsilon}^{1},\ldots,\mathcal{T}_{\varepsilon}^{\lambda}$, such that the collection of cells at each level in $\mathcal{T}$ is contained in exactly one of these $\varepsilon$-quadtrees, where $\lambda=\log_{2}\mathcal{E}$, see Eq. (2.1).
The problem with using a quadtree, for our purposes, is the alignment of the boundaries of their cells. Intuitively, two points that are close together might be separated by quadtree boundary, that belongs to cells that are dramatically bigger than this distance. Fortunately, this problem can be overcome by shifting (the point set, or the quadtree) $d+1$ times.
@@
Consider any two points $p,q\in[0,1)^{d}$, and let $\mathcal{T}$ be the infinite quadtree of $[0,2)^{d}$. For $D=2\left\lceil{d/2}\right\rceil$ and $i=0,\ldots,D$, let $\nu_{i}=(i/(D+1),\ldots,i/(D+1))$. Then there exists an $i\in\left\{{0,\ldots,D}\right\}$, such that $p+\nu_{i}$ and $q+\nu_{i}$ are contained in a cell of $\mathcal{T}$ with side length $\leq 2(D+1)\left\|{pq}\right\|$.
@@
The original proof is a direct consequence of a beautiful result of Walecki from the $19$th century [Als08], see [CHJ20] for details. For the sake of completeness, we sketch a shorter and inferior argument – a random permutation (i.e., order) of $S$ has a specific pair being consecutive with probability at least $1/\left|{S}\right|=1/t^{d}$. Repeating this process $m=O(t^{d}\log t)$ times, implies that the probability that there is any pair of elements in $S$ that is not consecutive, in one of the orderings, is $\leq\left|{S}\right|^{2}(1-1/t^{d})^{m}\ll 1/2$. Thus, the computed set of orderings has the desired property with constant probability.

### v3

@@ Abstract
For a parameter $\varepsilon\in(0,1)$, a set of orderings is $\varepsilon$-locality-sensitive (LSOs) if for any two points, $p,q\in[0,1)^{d}$, there exist an order in the set such that all the points between $p$ and $q$ (in the order) are $\varepsilon$-close to either $p$ or $q$. Since the original construction of LSOs can not be (significantly) improved, we present a construction of modified LSOs, that yields a smaller set, while preserving their usefulness. Specifically, the resulting set of LSOs has size $O(\mathcal{E}^{d-1}\log\mathcal{E})$, where $\mathcal{E}=1/\varepsilon$. This improves over previous work by a factor of $\mathcal{E}$, and is optimal up to a factor of $\log\mathcal{E}$.
This results in a flotilla of improved dynamic geometric algorithms, such as maintaining bichromatic closest pair, and spanners, among others. In particular, for geometric dynamic spanners the new result matches (up to the aforementioned $\log\mathcal{E}$ factor) the lower bound. Thus providing a near-optimal simple dynamic data-structure for maintaining spanners under insertions and deletions.
@@ 1 Introduction
Chan et al. chj-lota-20 showed that there is a “small” set $\Pi$ of orderings, such that for every pair of points $p,q\in\mathcal{H}$, there is an ordering $\sigma\in\Pi$ that is $\varepsilon$-local for them, where $\varepsilon\in(0,1)$ is a fixed parameter.
Definition 1.1.
For a pair of points $p,q\in\mathcal{H}=[0,1)^{d}$, an order $\sigma$ over the points of $\mathcal{H}$ is $\varepsilon$-local, for $p$ and $q$, if
$\sigma(p,q)\quad\subseteq\quad\mathcalb{b}\left({p,\varepsilon\ell}\right)\cup\mathcalb{b}\left({q,\varepsilon\ell}\right),$
where $\ell=\left\|{p-q}\right\|$, and $\mathcalb{b}\left({p,r}\right)$ denotes the ball of radius $r$ centered at $p$, see Figure 1.1 and Figure 1.2 (i).
Figure 1.1: All the points between $p$ and $q$ in $\sigma$ (i.e., the orange colored region) are inside the balls around $p$ and $q$.
Namely, all the points between $p$ and $q$ in $\sigma$ are in the vicinities of $p$ and $q$ in $\mathcal{H}$. Informally, there is an order in $\Pi$ that respects the localities of $p$ and $q$. Chan et al., inspired by locality sensitive hashing, referred to $\sigma$ as being a locality sensitive ordering (LSO).
Chan et al. showed that one can compute such a “universal” set of orderings $\Pi$, of size $O(\mathcal{E}^{d}\log\mathcal{E})$, where $\mathcal{E}=1/\varepsilon$. This set of orderings can be easily computed, and one can compute quickly the order of two points according to a specific order in the set. Having such a set of orderings reduces some dynamic geometric problems to the maintenance of a “few” sorted lists of points under insertions and deletions. Thus, certain problems in $d$ dimensions, are reduced to a collection of problems in one dimension. This can be interpreted as a somewhat strange dimension reduction result, (multi)embedding $[0,1)^{d}$ in the real line, such that for any two points in the hypercube, there is at least one embedding that has certain desired geometric property.
For example, using these orderings one can easily maintain a spanner under insertion/deletion operations with logarithmic time per operation (ignoring terms depending on $\varepsilon$). Specifically, one maintains (under updates) sorted lists of the point set, according to each of the orderings in $\Pi$. Here, an edge in the spanner is present only if the two points are adjacent in one of these ordered lists. Note that for each of these one dimensional instances, one maintains a dynamic ordered map of the points (e.g., balanced binary search tree).
The challenge..
So to reduce the number of orderings further, we have to relax the requirement somehow. A natural first step is to enlarge the vicinity to $p$ and $q$ to be an “$\varepsilon$-hippodrome” region.
Definition 1.2.
For a segment $pq$ in $\mathbb{R}^{d}$, and a number $r\geq 0$, let $\mathcalb{h}\left({pq,r}\right)$ be the hippodrome of $pq$ of radius $r$. Formally, it is the set of all the points in distance at most $r$ from $pq$: $\mathcalb{h}\left({pq,r}\right)=\left\{u\in\mathbb{R}^{d}\;\middle|\;\mathsf{d}\left({u,pq}\right)\leq r\right\},$ where $\mathsf{d}\left({u,pq}\right)=\min_{v\in pq}\left\|{u-v}\right\|$.
For a pair of points $p,q\in\mathcal{H}=[0,1)^{d}$, an order $\sigma$ over the points of $\mathcal{H}$ is $\varepsilon$-hippodrome local, for $p$ and $q$, if
$\sigma(p,q)\quad\subseteq\quad\mathcalb{h}\left({pq,\varepsilon\ell}\right),$
where $\ell=\left\|{p-q}\right\|$. See Figure 1.2 (i) and (ii).
@@
This $\varepsilon$-hippodrome locality is sufficient for some applications, but others require a stronger property of having a “gap” in the ordering.
@@
For $\gamma\in(0,1/4)$, with $\gamma\geq\varepsilon$ (e.g., $\gamma=1/4$), an order $\sigma$ is $(\varepsilon,\gamma)$-local, for $p$ and $q$, if
@@
$\sigma(p,q)\quad\subseteq\quad\mathcalb{h}\left({pq,\varepsilon\ell}\right)\cap\left({\Bigl{.}\mathcalb{b}\left({p,\gamma\ell}\right)\cup\mathcalb{b}\left({q,\gamma\ell}\right)}\right),\qquad\text{where}\qquad\ell=\left\|{p-q}\right\|,$
see Figure 1.2 (iii). Namely, the two sets $X=\sigma(p,q)\cap\mathcalb{h}\left({pq,\varepsilon\ell}\right)\cap\mathcalb{b}\left({p,\gamma\ell}\right)$ and $Y=\sigma(p,q)\cap\mathcalb{h}\left({pq,\varepsilon\ell}\right)\cap\mathcalb{b}\left({p,\gamma\ell}\right)$ have the gap property that $\mathsf{d}\left({X,Y}\right)\geq(1-2\gamma)\ell$.
Computation model..
@@
The model of computation used here is a unit-cost real RAM, supporting standard arithmetic operations and comparisons (but with the floor function), augmented with standard bitwise-logical operations (bitwise-exclusive-or and bitwise-and), which are widely available as assembly commands on all modern general purpose CPUs, and programming languages. This computational model is reasonable, and is used (for example) in working with compressed quadtrees efficiently h-gaa-11.
Previous related work..
@@
LSOs were introduced by Chan et al. chj-lota-20. Variations of LSOs for doubling metrics, and other metric spaces, were studied by Filtser and Le fl-loars-22 (Triangle LSO, Left-sided LSO), and Filtser f-label-23 (Triangle LSO, Rooted LSO). Since these works study LSO in the non-euclidean settings, they are not directly related to our work here (and they yield, naturally, worse bounds). LSOs were used to construct reliable spanners by Buchin et al. bho-sda-20, bho-srsal-22 in a simple plug & play fashion. There is some work on quadtrees for hyperbolic space by Kisfaludi-Bak and van Wordragen skg-quadt-24, which potentially can lead for LSOs for such spaces.
@@
We show that there is a universal set of LSOs of $[0,1)^{d}$ that is of size $O(\mathcal{E}^{d-1}\log\mathcal{E})$, such that these orderings have the somewhat weaker $(\varepsilon,\gamma)$-locality property, where $\gamma\geq\varepsilon$ is a large fixed constant. This improves the result of Chan et al. by a factor of $\mathcal{E}=1/\varepsilon$. Furthermore, for many of the applications using LSOs, one can replace the LSOs by the new set of “weaker” orderings, thus strictly improving their dependence on $\varepsilon$. In particular, for spanners the new construction enables maintaining dynamic spanners. Notably, up to $\log\mathcal{E}$ factor, the new construction matches the known lower bound on the number of edges in a spanner, in the static case, thus implying that the new orderings are almost optimal, and no approach based on LSOs can do (significantly) better in the Euclidean case.
Applications..
Since we have to slightly weaken the definition of LSOs to get the improved size, we have to rederive the proofs of correctness for some of the applications. LSOs are especially useful in applications where one wants to solve some geometric problem under insertions and deletions of points. Since LSOs decouple the geometric problem from the data-structure problem (i.e., all one needs to do is to maintain sorted lists of points under insertions and deletions), this results in simple data-structures that supports updates:
@@
    Bichromatic closest pair. In Section 4.2, we show how to use LSOs to dynamically maintain the bichromatic closest pair, as the shortest bichromatic edge in the locality graph is the desired approximation. This also readily leads to a data-structure that supports $(1+\varepsilon)$-approximate nearest-neighbor queries. We do not present the details for this here, as these results are, relatively speaking, less interesting.
@@
All these results improve by a factor of $\mathcal{E}$ over the previous results of Chan et al.. The new LSOs also improves, by a factor of $\mathcal{E}$, in a plug and play fashion other results, but the improvement is somewhat less interesting in these cases (i.e., the improvement is intuitively clear, and in some cases requires a lot of tedium). This includes reliable spanners bho-sda-20, dynamic approximate minimum spanning trees chj-lota-20, and vertex/edge fault-tolerant spanners chj-lota-20.
Sketch of the old construction..
We describe shortly the construction of LSOs by Chan et al. chj-lota-20, see Section 2.2 for more details. Chat et al. reduced the problem of computing LSOs to the following problem – consider the integer grid $B=\left\llbracket 4\mathcal{E}\right\rrbracket^{d}$, where $\mathcal{E}=1/\varepsilon$ and $\left\llbracket n\right\rrbracket=\{1,\ldots,n\}$. An edge of $G$ is long if it corresponds to a pair of points of $B$ that are in distance at least (say) $2/\varepsilon$ from each other. The long edges corresponds to well-separated pairs of points, and these are the ones we must “serve”. (Understanding why we concentrate on the long edges requires a bit deeper dive into the details, see Section 2.2.) To this end, consider the complete graph $G$ over $B$. Constructing LSOs is then achieved by constructing LSOs for $B$. A single ordering (i.e., LSO) for $B$ is simply a Hamiltonian path of $G$, and the task at hand is to compute a small number of Hamiltonian paths of $G$ that visits all the long edges. Specifically, a Hamiltonian path (in a graph) induces an ordering of the vertices, thus Hamiltonian paths serves as a way to encode the desired orderings. Having a long edge in such a path, ensures this pair is consecutive in the associated ordering.
Walecki a-wwc-08 showed a decomposition of the clique into Hamiltonian paths, and this leads to the desired LSOs of $B$. The problem with this approach is that it requires $\Omega(\mathcal{E}^{d})$ paths/LSOs. There is an additional blowup by a factor of $O(\log\mathcal{E})$ in the number of LSOs, when converting the LSOs of $B$ to LSOs of $[0,1)^{d}$.
Sketch of new approach..
@@
Why does the size of LSOs matter? LSOs seems to be algorithmically equivalent to other geometric tools, such as well-separated pair decomposition ck-dmpsa-95. Thus, they seem to be a fundamental property of the underlying space. It is thus natural to quantify exactly what the minimum size of LSOs needed for $\mathbb{R}^{d}$, as a function of $d$ and $\varepsilon$. Our work thus can be interpreted as (almost) settling this question.
@@
We describe the old construction of Chan et al. in Section 2.2. We present some necessary standard results on nets/packings in Section 2.3. In Section 3.1, we describe how to construct the needed packings on projected points, and how to decompose such a packing into a packing of larger radius. At last, we are ready. We present, in Section 3.2, the construction of the orderings with the hippodrome property, and, in Section 3.3, we present the modified construction with the gap property.
@@ 2 Preliminaries
2.1 Definitions
Definition 2.1.
For two sets $X,Y\subseteq\mathbb{R}^{d}$, let $\mathsf{d}\left({X,Y}\right)=\min_{x\in X,y\in Y}\left\|{x-y}\right\|$ denote the distance between $X$ and $Y$. For a point $p\in\mathbb{R}^{d}$, we use the shorthand $\mathsf{d}\left({p,X}\right)=\mathsf{d}\left({\{p\},X}\right)$.
@@ 2.2 Locality sensitive orderings
Our starting point is the result of Chan et al. chj-lota-20. Since we use the same framework, we quickly sketch the above construction (reproducing some relevant definitions).
@@
Informally, an $\varepsilon$-quadtree connects a node directly to its decedents that are $\lambda$ levels below it in the regular quadtree. Thus, to cover the quadtree, we need $\lambda$ different $\varepsilon$-quadtrees to cover all the levels of the quadtree. This can be achieved by setting directly the level of the root of the $\varepsilon$-quadtree. Specifically, we fix the root of the $\varepsilon$-quadtree $\mathcal{T}_{\varepsilon}^{i}$ to be $[0,2^{i})^{d}$, for $i=1,\ldots,\lambda$. This implies the following.
@@
Let $\mathcal{T}$ be a regular (infinite) quadtree over $[0,2)^{d}$. There are $\lambda$ $\varepsilon$-quadtrees $\mathcal{T}_{\varepsilon}^{1},\ldots,\mathcal{T}_{\varepsilon}^{\lambda}$, such that the collection of cells at each level in $\mathcal{T}$ is contained in exactly one of these $\varepsilon$-quadtrees, where $\lambda=\log_{2}\mathcal{E}$, see Eq. (2.1).
The problem with using a quadtree, for our purposes, is the alignment of the boundaries of their cells. Intuitively, two points that are close together might be separated by quadtree boundary, that belongs to cells that are dramatically bigger than the distance between the points. Fortunately, this problem can be overcome by shifting (the point set, or the quadtree) $d+1$ times.
@@
Consider any two points $p,q\in[0,1)^{d}$, and let $\mathcal{T}$ be the infinite quadtree of $[0,2)^{d}$. For $D=2\left\lceil{d/2}\right\rceil$ and $i=0,\ldots,D$, let $\nu_{i}=(i/(D+1),\ldots,i/(D+1))$. Then there exists an $i\in\left\{{0,\ldots,D}\right\}$, such that $p+\nu_{i}$ and $q+\nu_{i}$ are contained in a cell of $\mathcal{T}$ with side length $\leq 2(D+1)\left\|{p-q}\right\|$.
@@
The original proof is a direct consequence of a beautiful result of Walecki from the $19$th century a-wwc-08, see chj-lota-20 for details. For the sake of completeness, we sketch a shorter and inferior argument – a random permutation (i.e., order) of $S$ has a specific pair being consecutive with probability at least $1/\left|{S}\right|=1/t^{d}$. Repeating this process $m=O(t^{d}\log t)$ times, implies that the probability that there is any pair of elements in $S$ that is not consecutive, in one of the orderings, is $\leq\left|{S}\right|^{2}(1-1/t^{d})^{m}\ll 1/2$. Thus, the computed set of orderings has the desired property with constant probability.